OpenStudy (anonymous):

Need help on a question on Scientific notation. Question is in the comments.

sammixboo (sammixboo):

Question?

OpenStudy (anonymous):

What is the value of the expression, written in standard form? \[\left( 6.6x10^{-2} \right) divided by \left( 3.3x10^{-4} \right)\]

OpenStudy (anonymous):

@sammixboo

OpenStudy (unklerhaukus):

\[6.6\times10^{-2}\div3.3\times10^{-4}\\ =(6.6\div3.3)\times10^{-2-(-4)}\\ = ...\\ =...\times10^{...}\]

OpenStudy (anonymous):

Refer to the Mathematica attachment.

OpenStudy (anonymous):

http://gamutech.com We'll help and give you answers and there's a chat no lag like OS and you won't get suspended